OpenStudy (anonymous):

Suppose f(x)=4x-2 and g(x)=-2x+1. Find the value of f(5)/g(-3)

OpenStudy (anonymous):

f(5)=4(5)-2=20-2=18 g(-3)=-2(-3)+1=6+1=7 f(5)/g(-3)=18/7

OpenStudy (anonymous):

1 5/9 2 4/7 -2 2 are the choices

OpenStudy (anonymous):

mmmm

OpenStudy (anonymous):

i believe the answer is correct..??

zepdrix (zepdrix):

18/7? Hmm yah that sounds correct :) now change it to a mixed number so it matches one of your options.